[outages] Level 3 fiber cut San Bernardino, CA
ml at knight-networks.com
Wed Oct 24 08:43:01 EDT 2018
Got a Level 3 fiber cut in San Bernardino, CA going on for about 17
hours now. Our first link alarm was at 3:47pm CT yesterday. $DAYJOB
has two 1G circuits affected.
Looks like they had a hell of a time getting the replacement cable and
splice conduit installed. Splicing is estimated to start in another
2.5-3.5 hours on the east side of the cut.
Here's the latest we received from Level 3.
10/24/2018 12:29:56 GMT -
*** CASCADED EXTERNAL NOTES 24-Oct-2018 12:22:53 GMT From CASE: 15371425
The Outside Plant Engineer has reported that cable preparations on both
the west side and the east side of the cable is ongoing. It is estimated
splicing on the west side will begin at 14:00 GMT and splicing on the
east side will begin between 15:00 GMT and 16:00 GMT.
10/24/2018 11:10:34 GMT -
*** CASCADED EXTERNAL NOTES 24-Oct-2018 11:04:06 GMT From CASE: 15371425
The Outside Plant Engineer has confirmed the trench is dug and the
replacement cable is being ran between manholes at the west side of the
damage. Preparations are complete at the east side and splicing expected
to begin at 11:30 GMT with the 264 count cable, then the 84 count cable
followed by a 60 count lateral cable. Restoration efforts are ongoing
with an ETTR that remains unchanged, 17:30 GMT.
10/24/2018 10:05:30 GMT -
*** CASCADED EXTERNAL NOTES 24-Oct-2018 09:58:52 GMT From CASE: 15371425
The Outside Plant Engineer reported the replacement section of cable got
stuck when pulling it into position. A second splice enclosure will be
installed in the next manhole. Field Operations will dig a trench
between the two enclosures to complete the installation of the 2,400
foot section of 264 count replacement cable. The trench is expected to
be complete at approximately 10:30 GMT with an estimated time splicing
will begin of 12:30 GMT.
10/24/2018 08:08:37 GMT -
*** CASCADED EXTERNAL NOTES 24-Oct-2018 08:01:50 GMT From CASE: 15371425
The Outside Plant Engineer reported the replacement section of fiber
cable has arrived and is currently being pulled into position. Along
with the section of fiber cable, a new splice enclosure is being prepped
and will be replaced to complete repairs of the 264 count cable. The 84
count cable has been prepped and ready for splicing. Field Operations
have moved to an adjacent manhole looking for available slack to repair
to the 168 count cable. The Outside Plant Engineer reported splicing is
expected to commence at 10:00 GMT. The next update will be sent at that
time or if additional information is received.
More information about the Outages